# Higher-Order Equivariant Neural Networks for Charge Density Prediction in Materials

### v1

@@ 2.3 Effects of higher-order representations
To demonstrate the impact of introducing higher-order equivariant tensor representations in our model, we examine the effect of training models while varying the maximum rotation order $L$, with $\ell\in\{0,...,L\}$ up to $L=4$. While higher-order representations are achievable, they can be prohibitive to use in the network due to the $O(L^{6})$ computational complexity of tensor products [26]. We determine the number of channels for each rotation order by $N_{\text{channels}}=\lfloor\frac{500/(L+1)}{\ell*2+1}\rfloor$. For example, the $L=0$ model has representations consisting of 500 even scalars, and 500 odd scalars, while the $L=1$ model has representations consisting of 250 even scalars, 250 odd scalars, 83 even vectors, and 83 odd vectors. In this way, each model is constructed such that the total representation size and the proportion of the total representation used by each rotation order are approximately equal. We train each model on 1,000, and 10,000 material subsets and the full MP dataset, using the same validation and test sets. As shown in Fig. A.1, we observe a consistent increase in performance on each training subset as the maximum rotation order increases, and a similar scaling behavior with respect to the training set size. This trend suggests that higher-order representations are necessary for accurate charge density modeling, and can match the performance of lower order models with less data.
@@
At the atomic level, we find more evidence of the benefits of higher-order equivariant representations in modeling charge density. We examine the per-element errors by computing $\epsilon_{\text{mae}}$ on grid points $G_{\text{local},i}\subset G$, a set of points that are within 2Å of an atom of species $i$. Figure 3 shows this per-element error, averaged across the test set. While the $L=0$ model struggles with reactive non-metals and metalloids, which often form covalent bonds, $L=4$ shows strong improvement among these elements. Nearly all elements show improvement, but the median change in $\epsilon_{mae}$ values for those materials containing a non-metal or metalloid is $-44.6\%$ ($n=1628$) compared to $-23.0\%$ for materials with only metals ($n=346$) (Fig. A.2).
Figure 3: Effect of rotation order on element-wise error. Average $\epsilon_{\text{mae}}$ per element at $L=0$ (left) and $L=4$ (right), showing performance improvements across the periodic table. Values are computed over the set of grid points $G_{\text{local}}$, and averaged across all materials containing those elements in the test set. Only elements in at least 10 test set materials are shown.
@@
where $G$ is a set of probe points and $\hat{r}_{ki}$ is a unit vector from probe point at location $\vec{g}_{k}$ to the closest atom at location $\vec{r}_{i}$. Intuitively, a material with charge density gradients pointing directly towards or away from the nearest atom will have dot products larger in magnitude and $\zeta\rightarrow 0$, whereas a material with density that has high angular variance with respect to the nearest atom will have dot products smaller in magnitude, and $\zeta\rightarrow 1$. Figure A.3 shows the test set average $\zeta$ over $G_{local,i}$ for every element. Figure A.4 demonstrates that the differences in performance between the lower and higher rotation order networks correlates strongly with $\zeta$. With covalent bonding and associated high angular variance frequently present in the Materials Project data, this further justifies the need for introducing higher rotation order representations into charge density prediction networks.
@@ 2.4 Runtime and scalability
Figure 4: Runtime and acceleration of DFT. a: Runtime comparison of DFT and our ChargE3Net model with respect to number of atoms in the evaluated system. DFT is using a 48-core Intel Xeon CPU, while ChargE3Net uses a single NVIDIA V100 GPU. b: Number of SCF steps required to reach an energy change threshold of $10^{-5}$ for non-magnetic materials in the test set. Box plot edges show 25th and 75th percentiles, while the white line in the box plot shows the median. SC = self-consistent calculations starting from SAD. SC from ChargE3Net = self-consistent calculations starting from a charge density estimate with ChargE3Net. SC from SC = self-consistent calculation using a charge density initialized from the SC calculation of the same material. NSC from $x$ = non-self-consistent calculation using the $x$ starting point.
2.5 Using predicted charge densities to accelerate DFT
DFT calculations require an initial charge density and wavefunction prior to iteratively solving the Kohn-Sham equations until self-consistency [27]. Charge density initializations that closely match the ground-state charge density often save compute time by reducing the number of self-consistent field (SCF) steps needed to find a self-consistent solution. Many schemes for initializing the charge density exist in the literature, but perhaps the most common one is the superposition of atomic densities (SAD), where the initial charge density is computed by summing over the atomic charge density of each atom in the system [28].
We show that charge densities predicted by ChargE3Net reduce the number of SCF steps, thereby accelerating DFT calculations, on unseen non-magnetic materials (see Fig. 4b) compared to a SAD initialization. The median number of SCF steps needed to converge a single-point calculation self-consistently for a non-magnetic material starting from the SAD initialization is 15, while the same starting from a ChargE3Net initialization is 11, representing a 26.7% decrease in SCF steps. As a benchmark, we also include the distribution of the number of SCF steps needed to converge a calculation starting from a converged (ie ground-truth) charge density (“SC from SC”), showing that our model still has room for improvement. About 96% of calculations done with a predicted charge density ran with fewer SCF steps than initializing with SAD, and the compute time improvement is robust to a wide range of $\epsilon_{mae}$ (see Fig. A.7).
Figure 5: Non-self-consistent physical property calculation. Absolute errors of the total energy normalized by the number of atoms (top right), atomic forces (top left), band energies (bottom right), and band gaps (bottom left). Dashed lines indicate chemical accuracy (43 meV) except for per-atom energies, which is 1 meV/atom, and forces, which is 0.03 eV/Å.
In Fig. 5 we compute the energy, forces, band energies, and bandgaps of the test set non-self-consistently to evaluate how well ChargE3Net’s predicted charge densities reproduce physical properties. 39% of materials in the test set have per-atom energy errors less than 1 meV/atom. 67% of materials have forces less than 0.03 eV/Å, a common threshold for structural relaxations; 15% of the test set have force errors of 0.0. All of the structures with perfect agreement have self-consistent forces of 0.0 eV/Å for each atom and Cartesian component, possibly because of cancellation of charge density errors due to symmetry. 43% and 75% of materials recover band energies at each k-point and the direct band gap within chemical accuracy (43 meV), respectively. 45% of the band gaps obtained non-self-consistently exactly matched those of self-consistent calculations. Similar to the force data described above, calculations with perfect agreement all have a self-consistent band gap of 0 eV, indicating the model is able to readily identify metallic materials. We note that our use of chemical accuracy for band gaps is a more stringent criterion than experimental measurement errors for well-studied semiconductors [30].
In addition to the model’s robustness to materials outside its training set, we show that ChargE3Net can predict the charge densities of materials after small atomic perturbations with minor increase to the charge density error (see Fig. A.6). This allows us to accurately obtain phonon properties, which may rely on finite differences as the main method to obtain approximations to the local potential energy surface, from non-self-consistent calculations. We compute the constant-volume heat capacity ($C_{v}$) under the harmonic approximation for two zeolites, which are not included in the Materials Project dataset, and show that $C_{v}(T)$ obtained closely match those obtained by self-consistent calculations (see Section A.1). As another example of transferability, we compute the electronic bandstructure of the Si(111)-2x1 surface non-self-consistently from the ChargE3Net prediction and show that it almost exactly matches that of a self-consistent calculation (see Section A.2).
@@ 3 Discussion
This work introduces ChargE3Net, an architecture for predicting charge density through equivariance with higher-order representations. We demonstrate that introducing higher-order tensor representations over scalars and vectors used in prior work achieves greater predictive accuracy, and show how this achieved through the improved modelling of systems with high angular variance. ChargE3Net enables accurate charge density calculations on atomic systems larger than what is computationally feasible with DFT, but can also be used in combination with self-consistent DFT to reduce the number of SCF steps needed for convergence. Our findings hold across the large and diverse Materials Project dataset, paving the way towards a universal charge density model for materials.
@@
The work of Hohenberg and Kohn proves that all material properties could, in principle, be derived from the electron density [1]. Because this quantity is so fundamental to predicting material properties, we view models that compute electron densities as a potential intermediate towards other material properties that may improve the generalizability of machine learning models in material science. This leads us to believe that large models trained on a dataset that spans the known material space could be a promising direction towards a foundation model for DFT calculations that can be fine-tuned for additional properties. As evidence, our non-self-consistent calculations, which can be seen as a ground-truth calculation of the electronic strucuture (and derived properties) using the crystal structure and charge density as inputs, demonstrate that one could indeed obtain electronic and thermal properties of materials using highly accurate charge-density estimates (see Sections A.1 and A.2). On a broader dataset, we show that non-self-consistent calculations are able to capture the total energy, atomic forces, electronic bandstructure, and the band gap within chemical accuracy for a majority of structures in the test set, though there is room for improvement. We show in Fig. A.8 - A.13 that errors in the charge density establish a lower bound to the property error, and that small charge density errors can give rise to substantial property error. We hypothesize that small errors localized around atomic nuclei could be the reason for higher than expected property errors, though further study is needed to confirm this.
@@
While higher order SO(3) equivariant tensor operations yield significantly improvement in charge density predictions (Fig. 2 and 3), the SO(3) convolution used in this work are limited by the $O(L^{6})$ time complexity of tensor products. Recent work [26] has demonstrated that SO(3) convolutions can be reduced to SO(2), with a time complexity of $O(L^{3})$. In the future, our approach can be combined with the methodology presented in [26] to improve computational efficiency, and enable the practical usage of higher rotation orders ($L>>4$) in atomistic modeling.
@@ 4 Methods
4.1 Architecture
$A^{n}$ is initialized as a one-hot encoded $z$, represented as a $\ell=0$, $p=1$ tensor with $N_{\text{channels}}$ equal to the number of atomic species. $P^{n}$ is initialized as single scalar zero, with $\ell=0$, $p=1$, and $N_{\text{channels}}=1$. Each representation is updated through a series of alternating convolution $\text{Conv}(\cdot)$ and non-linearity $\text{Gate}(\cdot)$ layers, a total of $N_{\text{interactions}}=3$ times. With each convolution layer, all possible outputs of the tensor product at different $\ell$ and $p$ values with $\ell_{o}\leq L$ are concatenated, increasing the representation size past the scalar initializations of $A^{n}$ and $P^{n}$. Atom representations are updated with $A_{i}^{n+1}=\text{Gate}(\text{Conv}^{n}_{\text{atom}}(\vec{r}_{i},A_{i}^{n}))$. $\text{Conv}_{\text{atom}}$ is defined as:
@@ 4.1 Architecture
4.2 Training details
4.2.1 Optimization
Table A.1 outlines the experimental setup for training ChargE3Net on each of the datasets. All models For each gradient step, a random batch of materials is selected, from which a random subset of the charge density probe points are used. We use the Adam optimizer [35], and decay the learning rate by $0.96^{s/\beta}$ at step $s$. We find that optimizing for L1 error improves performance and training stability over mean squared error (see Table A.2).
4.2.2 Graph construction
We propose an algorithm to achieve more efficient atom-probe graph construction, enabling faster training and inference times. Prior work constructs graphs by adding probe points directly to the atomic structure, computing edges with atomic simulation software such as ase or asap [36], then removing edges between probes. While this is performant for small systems with few probes, we find it does not scale well due to the redundant computation of probe-probe edges, which grow quadratically with respect to the number of probe points. To address this, our algorithm leverages two k-d trees [37]; one to partition atoms, and one to partition probe points. This enables an efficient radius search for atom-probe edges exclusively. We implement this through modification of the k-d tree implementation scipy [38] to handle periodic boundary conditions. Constructing graphs with 1000 probe points, we observe an average graph construction time ($\pm$ standard error) of $0.00758\pm 0.00087$, $0.0103\pm 0.0005$, and $0.00462\pm 0.00056$ seconds for MP, QM9, and NMC respectively. This is at least an order of magnitude faster than methods from [11], with which we measured graph construction times of $12.6\pm 0.9$, $0.114\pm 0.006$, and $0.141\pm 0.010$ seconds.
@@ 4.3 Datasets
The QM9 and NMC datasets, along with train, validation and test splits, are provided by Jørgensen and Bhowmik [11]. The QM9 dataset[23] contains VASP calculations of the 133,845 small organic molecules introduced by Ramakrishnan et al. [22], with training, validation, and test splits of size 123835, 50, and 10,000 respectively. The NMC dataset[24] consists of VASP calculations of 2000 randomly sampled nickel manganese cobalt oxides containing varying levels of lithium content, with training, validation, and test splits of size 1450, 50, and 500 respectively. For the MP data, we collected 122,689 structures and associated charge densities from api.materialsproject.org. Structures in the dataset that shared composition and space group were identified as duplicates and only the highest material_id structure was included, leaving 108,683 materials. The data was split into training, validation, and test splits with sizes 106,171, 512, and 2000 respectively. Later, analyzing the test set materials for thermodynamic stability and structural abnormalities (Fig. A.5), 26 materials were found to have unrealistic volume per atom, in excess of $100$${}^{3}$/atom and mostly from space group Immm. These materials were removed from the test set and excluded from results.

### v2

@@ 2.4 Higher-order representations lead to improved model accuracy
To demonstrate the impact of introducing higher-order equivariant tensor representations in our model, we examine the effect of training models while varying the maximum rotation order $L$, with $\ell\in\{0,...,L\}$ up to $L=4$ used in the final model. We determine the number of channels for each rotation order by $N_{\text{channels}}=\lfloor\frac{500/(L+1)}{\ell*2+1}\rfloor$. For example, the $L=0$ model has representations consisting of 500 even scalars, and 500 odd scalars, while the $L=1$ model has representations consisting of 250 even scalars, 250 odd scalars, 83 even vectors, and 83 odd vectors. In this way, each model is constructed such that the total representation size and the proportion of the total representation used by each rotation order are approximately equal. We train each model on 1,000, and 10,000 material subsets and the full MP dataset, using the same validation and test sets. We observe a consistent increase in performance on each training subset as the maximum rotation order increases, and a similar scaling behavior with respect to the training set size (Supplementary Fig. A.12). This trend suggests that higher-order representations are necessary for accurate charge density modeling, and can match the performance of lower order models with less data.
@@
At the atomic level, we find more evidence of the benefits of higher-order equivariant representations in modeling charge density. We examine the per-element errors by computing $\epsilon_{\text{mae}}$ on grid points $G_{\text{local},i}\subset G$, a set of points that are within 2Å of an atom of species $i$. Fig. 4 shows this per-element error for both $L=0$ (lower-left triangles) and $L=4$ (upper-right triangles), averaged across the test set. While the $L=0$ model struggles with reactive non-metals and metalloids, which often form covalent bonds, $L=4$ shows strong improvement among these elements. Nearly all elements show improvement, but the median change in $\epsilon_{mae}$ values for those materials containing a non-metal or metalloid is $-44.6\%$ ($n=1628$) compared to $-23.0\%$ for materials with only metals ($n=346$) (Supplementary Fig. A.13).
Figure 4: Effect of rotation order on element-wise error. The lower-left triangles indicate the average $\epsilon_{\text{mae}}$ per element at $L=0$, and the upper right triangles represent $\epsilon_{\text{mae}}$ per element at $L=4$. Performance improvements are shown across the periodic table. Values are computed over the set of grid points $G_{\text{local}}$, and averaged across all materials containing those elements in the test set. Only elements in at least 10 test set materials are shown.
@@
where $G$ is a set of probe points and $\hat{r}_{ki}$ is a unit vector from probe point at location $\vec{g}_{k}$ to the closest atom at location $\vec{r}_{i}$. Intuitively, a material with charge density gradients pointing directly towards or away from the nearest atom will have dot products larger in magnitude and $\zeta\rightarrow 0$, whereas a material with density that has high angular variance with respect to the nearest atom will have dot products smaller in magnitude, and $\zeta\rightarrow 1$; see Supplementary Fig. A.14 for the test set average $\zeta$ over $G_{local,i}$ for every element. We observe that the differences in performance between the lower and higher rotation order networks correlates strongly with $\zeta$ (Supplementary Fig. A.15). With covalent bonding and associated high angular variance frequently present in the Materials Project data, this further justifies the need for introducing higher rotation order representations into charge density prediction networks. In addition, we find the higher-order model maintains higher accuracy under perturbations to atomic positions (Supplementary Fig. A.11), an imperative trait when using models for phonon calculations.
@@ 3 Discussion
This work introduces ChargE3Net, an architecture for predicting charge density through equivariance with higher-order representations. We demonstrate that introducing higher-order tensor representations over scalars and vectors used in prior work achieves greater predictive accuracy, and show how this is achieved through the improved modelling of systems with high angular variance. ChargE3Net enables accurate charge density calculations on atomic systems larger than what is computationally feasible with DFT, but can also be used in combination with self-consistent DFT to reduce the number of SCF steps needed for convergence. Our findings hold across the large and diverse Materials Project dataset, paving the way towards a universal charge density model for materials. Future work to expand the model to charge density calculations of unrelaxed structures such as those in MPtrj [24] would support general application of ML models to accelerate ab initio calculations and should be pursued.
@@
The work of Hohenberg and Kohn proves that all material properties could, in principle, be derived from the electron density [1]. Because this quantity is so fundamental to predicting material properties, we view models that compute electron densities as a potential intermediate towards other material properties that may improve the generalizability of machine learning models in material science. As evidence, our non-self-consistent calculations, which can be seen as a ground-truth calculation of the electronic strucuture (and derived properties) using the crystal structure and charge density as inputs, demonstrate that one could indeed obtain electronic and thermal properties of materials using highly accurate charge-density estimates. On a broader dataset, we show that non-self-consistent calculations are able to capture the total energy, atomic forces, electronic bandstructure, and the band gap within chemical accuracy for a majority of structures in the test set, though there is room for improvement. This leads us to believe that large models trained on a charge density dataset that spans the known material space could be a promising direction towards a foundation model for DFT calculations that can be fine-tuned for downstream property prediction.
We have also explored the connection between errors in charge density predictions and the property error. We show in Supplementary Fig. A.17 - A.22 that errors in the charge density establish a lower bound to the property error, and that small charge density errors can give rise to substantial property error. Examining charge density errors more closely, specifically analyzing $\epsilon_{mae}$ localization in low- and high-error non-self-consistent calculations, we find difficulty identifying localized regions where charge density prediction errors lead to property errors (Supplementary Fig. A.23). While the errors in non-self-consistent property predictions are certainly affected by errors in charge density predictions from ChargE3Net, the errors may in part be explained by an incorrect DFT setting (the VASP tag LMAXMIX. See the Methods section for more information). The errors introduced by this calculation parameter may mean achieving chemical accuracy is improbable for some materials using non-self-consistent calculations even when initializing with a self-consistent (i.e. ground truth) charge density.
@@
We demonstrate higher-order E(3) equivariant tensor operations yield significant improvement in charge density predictions and are more data efficient (Fig. 3 and 4). However, the E(3) convolution used in this work are limited by the $O(L^{6})$ time complexity of tensor products [62]. Recent work [62] has showed that SO(3) convolutions can be reduced to SO(2), with a time complexity of $O(L^{3})$. In the future, our approach can be combined with the methodology presented in [62] to improve computational efficiency, and enable the practical usage of higher rotation orders ($L>>4$) in atomistic modeling.
@@ 4 Methods
4.1 ChargE3Net architecture
As described in Section 2.1, the inputs of the model consist of probe coordinates $\vec{g}$ and atom coordinates $\vec{r}$ with species $z$. Each atom and probe point are represented by a dictionary of tensors, where $A^{n}$ denotes an atom representation and $P^{n}$ denotes a probe representation, each at layer $n$. $A^{n=0}$ is initialized as a one-hot encoded $z$, represented as a $\ell=0$, $p=1$ tensor with $N_{\text{channels}}$ equal to the number of atomic species. $P^{n=0}$ is initialized as single scalar zero, with $\ell=0$, $p=1$, and $N_{\text{channels}}=1$. Each representation is updated through a series of alternating convolution $\text{Conv}(\cdot)$ and non-linearity $\text{Gate}(\cdot)$ layers, a total of $N_{\text{interactions}}=3$ times. With each convolution layer, all possible outputs of the tensor product at different $\ell$ and $p$ values with $\ell_{o}\leq L$ are concatenated, increasing the representation size past the scalar initializations of $A^{n}$ and $P^{n}$. Atom representations are updated with $A_{i}^{n+1}=\text{Gate}(\text{Conv}^{n}_{\text{atom}}(\vec{r}_{i},A_{i}^{n}))$. $\text{Conv}_{\text{atom}}$ is defined as:
@@ 4.1 ChargE3Net architecture
4.2 Training ChargE3Net models
We develop an algorithm to achieve more efficient atom-probe graph construction, enabling faster training and inference times. Prior work constructs graphs by adding probe points directly to the atomic structure, computing edges with atomic simulation software such as ase or asap [67], then removing edges between probes. While this is performant for small systems with few probes, we find it does not scale well due to the redundant computation of probe-probe edges, which grow quadratically with respect to the number of probe points. To address this, our algorithm leverages two k-d trees [68]; one to partition atoms, and one to partition probe points. This enables an efficient radius search for atom-probe edges exclusively. We implement this through modification of the k-d tree implementation scipy [69] to handle periodic boundary conditions. Constructing graphs with 1000 probe points, we observe an average graph construction time ($\pm$ standard error) of $0.00758\pm 0.00087$, $0.0103\pm 0.0005$, and $0.00462\pm 0.00056$ seconds for MP, QM9, and NMC respectively. This is at least an order of magnitude faster than methods from [43], with which we measured graph construction times of $12.6\pm 0.9$, $0.114\pm 0.006$, and $0.141\pm 0.010$ seconds.
To train the ChargE3Net model, Supplementary Table A.1 outlines the experimental setup on each of the datasets (e.g., learning rate, batch size, number of probe points and number of training steps). Specifically, at each gradient step, a random batch of materials is selected, from which a random subset of $N_{\text{points}}$ probe points are selected. We use the Adam optimizer [70], and decay the learning rate by $0.96^{s/\beta}$ at step $s$. We find that optimizing for L1 error improves performance and training stability over mean squared error (Supplementary Table A.2). Supplementary Fig. A.24 shows the evolution of training loss and validation error as the model is trained on each dataset. All models were implemented in PyTorch and trained on NVIDIA Volta V100 GPUs using a distributed compute architecture.
@@ 4.3 Datasets
The QM9 and NMC datasets, along with train, validation and test splits, are provided by Jørgensen and Bhowmik [43]. The QM9 dataset[56] contains VASP calculations of the 133,845 small organic molecules introduced by Ramakrishnan et al. [55], with training, validation, and test splits of size 123835, 50, and 10,000 respectively. The NMC dataset[57] consists of VASP calculations of 2000 randomly sampled nickel manganese cobalt oxides containing varying levels of lithium content, with training, validation, and test splits of size 1450, 50, and 500 respectively. For the MP data, we collected 122,689 structures and associated charge densities from api.materialsproject.org [49]. Structures in the dataset that shared composition and space group were identified as duplicates and only the highest material_id structure was included, leaving 108,683 materials. The data was split randomly into training, validation, and test splits with sizes 106,171, 512, and 2000 respectively. Later, analyzing the test set materials for thermodynamic stability and structural abnormalities (Supplementary Fig. A.25), 26 materials were found to have unrealistic volume per atom, in excess of $100$3/atom and mostly from space group Immm. These materials were removed from the test set and excluded from all reported results and figures.
We note to the reader a known issue with the Materials Project dataset. Original calculation parameters may have an incorrectly set VASP parameter called LMAXMIX, which sets the maximum angular momentum number for the one-center PAW charge densities saved in the charge density output of VASP and passed into the charge density mixer. This affects elements with d- and f-electrons when no Hubbard U parameter was set by the Materials Project. In Supplementary Fig. A.7, we correct this error for the randomly sampled test set, then compared the distribution of properties between non-self-consistent calculations of the original and corrected input parameters. Note that non-self-consistent calculations with a potentially incorrectly set LMAXMIX, as is set in the current Materials Project charge density dataset, makes a per-atom energy error about two orders of magnitude larger than that of self-consistent calculations. This represents an additional source of error to all non-self-consistent calculations based on Materials Project data.
In addition to a held-out test set randomly sampled from the Materials Project, we obtain a random set of 2000 published structures from GNoME [50] as a second test set. Specifically, we choose a random set of materials from GNoME with 5 or 6 unique species as these are less well-represented in the Materials Project training data. DFT calculations were set up using pymatgen [71] and otherwise using the same methodology as detailed in Section 4.4. 76 calculations failed, leaving 1924 materials for analysis, 1510 of which were non-magnetic. Non-magnetic structures were identified via DFT calculations initialized by SAD and magnetic moment guesses provided by pymatgen with an overall absolute magnetic moment <0.1 $\mu B$ and absolute atomic magnetic moments <0.1 $\mu B$ for every atom in the structure.
Supplementary Fig. A.1 shows a visualization of SOAP descriptors of materials in the MP train and test set and GNoME subset, projected into 2-d using t-SNE. We do not find a significant difference in distributions across the MP train and test sets, which is to be expected as the data is divided randomly. Figure A.2 shows the distribution of crystal systems and space groups within the MP training set. As seen with the SOAP descriptors, in Fig. A.3 we note a similar distribution of crystal systems across the MP train and test sets, with the GNoME subset showing significantly more monoclinic systems and significantly fewer tetragonal and cubic systems.